Let ? be the binomial random variable the outcome of performing 10 Bernoulli trials with the probability of success ? = 0.45. Suppose the outcome is ? = 1, find the probability that the only success has occurred at ??h trial.
Answer #1

Solution

Given that ,

p = 0.45

1 - p = 1 - 0.45 = 0.55

n = 10

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X = 1) = (10 C 1) * (0.45)1 * (0.55)9

= 0.020724

Probability = 0.0207